If a simple graph G has p vertices and any two distinct vertices u and v of G have the

property that degGu + degGv ≥ p-1 then prove that G is connected


1
Expert's answer
2021-08-11T05:09:36-0400


Assume that "G" is disconnected, that is there exist two vertices, "u" and "v", such that there is no path in "G" between them. Let "U" be the set of vertices incident to the vertex "u", and "V" be the set of vertices incident to the vertex "v". Since "u" and "v" are not incident, then "u\notin V" and "v\notin U".

If "U\cap V=\empty" then all the sets "\{u\},\,\{v\},\,U,\,V" are disjoint.

Hence, "|\{u\}|+|\{v\}|+|U|+|V|\leq |G|", or equivalently, "1+1+\deg_Gu+\deg_Gv\leq p", and "\deg_Gu+\deg_Gv\leq p-2", that is a contradiction to the condition "\deg_Gu+\deg_Gv\geq p-1".

From this it follows that the assumption "U\cap V=\empty" is false, i.e. there exists a vertex "w\in U\cap V". Thus, "w" is incident to both the vertices "u" and "v", and "u-w-v" is a path in "G", connecting "u" and "v", that is a contradiction with assumption that "G" is disconnected.

Therefore, "G" is a connected graph.